You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100615 (16 Jun 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 June14-15

Observer: V. Testa, K.Boutsia
Telescope Operator: D.Gonzales Huerta
Support Astronomer: O. Kuhn
Instrument: LUCIFER1 - LBC

Summary

Candidate photometric - seeing <1.0
Problems with the telemetry of the Blue channel. Last part of the night we switched to Red Only mode.
Time lost due to telemetry errors : >2hrs for Blue channel and 0.5h for Red.

Data Summary

Project Objects Filter ExpTime Notes
Mannucci HDFN 210zJHK H 2hr arc- flat- tell
Giallongo Q1623 Uspec, z ~40min blu errors during 2nd OB
Piotto NGC 6791 V-BESSEL, i-SLOAN 3000s each monocular only red, 2.2 h total obs. time

Details

02:21 - Opening the dome - starting interface

02:30 - test Flats to monitor sky brightness 0001-0060

02:35 - Sky Flat fields

in the K band - imaging
  • 2secs / 11000-9000counts 0001-0010
  • 2secs / 7000-6000counts 0011-0020
  • 2secs / 4500-4000counts 0021-0030

in the H band
  • 2secs / 9000-7000counts 0031-0040
  • 2secs / 6000-4000counts 0041-0050
  • 2secs / <3000counts 0051-0060

T/O corrected pointing from his position. Double check with pointing star

03:00 - Pointing Star near the field (WT10_385) 0061-0062
  • ptcorr, vey small correcton, procedure worked fine

03:06 - Persson standard BS9145 for collimation (POSITION,ACTIVE)

03:16 - Collimation - preset to Mannucci HDFN acquisition script for MOS
  • Lost guiding star, trying alternate one. No way. Check pointing again.

03:25 - Pointing ok, maybe sky still too bright.

03:31 - Change guide star, ok. Collimating

03:37 - MOS acquistion on HDFN K 0063-0067

03:45 - only 4 stars for reference 2,3,4,7 skipped
  • do a through-slit again to check
  • +0.12 in X to better center the ref. star into the slit

03:58 - HDFN H mos science script 3
  • 300DIT x 1NDIT x 12 exp 0068-0079

Give Y+1.5 offset to go back to original pointing

05:07 - through slit (0080-0086 discard) 0087

05:13 - HDFN H mos science script 4
  • 300 DIT x 1NDIT x 12 exp 0088-0099

06:20 - Calibration unit in: arcs and flats
  • ARCS Ar OFF 0100
  • ARCS Ar ON 0101
  • ARCS Xe OFF 0102
  • ARCS Xe ON 0103
  • FLAT Halo1 OFF 8sec 0104-0106
  • FLAT Halo1 ON 8sec 0107-0109
  • We have doubled the exposure time for the flat aquisition

06:36 - Telluric 0111-0150
  • not integrated but normal, every DIT saved separately
  • exposure time 3DITx10NDIT

06:41 - SWITCH TO LBC

07:01 - Switch complete
  • Go checking the poiting WT10_298
  • ptautoadjust ptmodify IE -22.153254121718 ptmodify CA -66.44293373613

07:11 - Slew to Q1623
  • dofpia
  • RED 4194 780 12 -1533 495 127 -14 0.85 1.82
  • BLUE -14720 0 0 0 0 0 0 0.43 5.80
  • RED -571 -235 -41 142 -2 290 25 0.82 0.32
  • BLUE -963 -1148 -827 -894 -188 -628 34 0.94 0.81
  • RED 723 -119 185 -71 77 -218 13 0.80 0.34
  • BLUE -275 -187 94 -114 -182 -254 -4 1.11 0.24
  • RED -329 -242 -7 83 -118 60 13 0.76 0.18 (come on!!!! Z4=-329 and analyzing !!! 250 as limit is ridiculous!)
  • BLUE -68 -17 100 106 90 -105 0 1.05 0.13
  • RED -154 210 54 54 19 -53 8 0.78 0.13
  • BLUE -162 -88 -248 -119 -38 38 -13 1.14 0.14

07:19 - Q1623_Uspec_2
  • in Uspec does not do IA online, probably no objects in the tech chip

07:57 - dofpia
  • RED -786 -47 72 200 -212 145 3 0.77 0.36
  • BLUE 264 789 59 228 204 -423 -29 1.03 0.40
  • RED 165 -64 -183 82 -5 -127 -18 0.61 0.15
  • BLUE -191 198 199 -63 -8 -177 -17 0.90 0.18

(telemetry error in the blue temperature, gave temperature limit exceeded 17.4 instead of -50 celsius)

08:02 - Q1623_Uspec_3 *08:09 - again temperature liits exceeded spoiled 5 min exposure).... 08:21 - again!!!!!! f!@@#@#$@ telemetry!!!!! CCD controller error, OB exited in error *08:26 - Power OFF/ Power ON then move to 6791 red only

08:35 - Slew to NGC 6791 try binocular first
  • dofpia
  • RED 641 -70 273 3 -208 -477 -45 0.49 0.42
  • BLUE 3798 -303 360 -39 -667 -781 -83 0.68 1.61
  • RED -718 38 -490 88 -9 89 -2 0.45 0.33
  • BLUE -371 -165 -337 44 277 -234 -20 0.61 0.28
  • RED 892 -292 -156 -141 -33 -183 -1 0.44 0.40
  • BLUE -317 -172 425 -180 114 59 -15 0.55 0.23
  • RED 8 475 211 -9 37 -13 -20 0.53 0.17
  • BLUE -344 364 -572 -41 51 125 -14 0.52 0.27
  • too many iterations, blue ccd controller error. Switch off blue, go on with red only.

08:50 - Red Only, dofpia (strangely, it did not took the fastextra OB, but ran the WD6791 that had been stopped while pointing....)
  • RED 1242 -171 -49 35 -158 -109 11 0.73 0.51
  • RED -73 396 9 182 36 0 0 0.62 0.17
  • RED -284 111 -119 -25 -149 111 17 0.74 0.17
  • RED 181 -133 100 -38 -9 -13 1 0.61 0.12

09:02 - WD_N671_LONG_01 Only red

10:06 - dofpia
  • RED -955 1865 2785 754 273 -2 64 0.79 1.12
  • RED 2007 -1539 -5161 -2646 -239 -452 37 0.00 2.15
  • RED -792 -1527 1706 633 -173 -713 -100 0.43 0.92
  • RED -1154 -64 201 88 57 144 -34 0.53 0.48
  • RED -223 116 -17 -16 4 90 -10 0.61 0.14

10:17 - WD_N671_LONG_02 Only red * Sky brightens at the end, twilight ongoing * last two skipped for saturation

11:20 - SkyFlatTes_redV
  • twice , second one has ~1000 cts/s

11:26 - SkyFlat_redVi_5 V=5s, i=10s
  • after the 8th, reduced to 3,6

11:32 - Again but changing dither pattern and V=1s, i=2s

11:38 - Again but changing dither pattern and V=1s, i=1s

Meanwhile sent DoDarks for Lucifer

11:45 - End of Night

-- VincenzoTest - 15 Jun 2010
Topic revision: r6 - 16 Jun 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback